I need a one example of how to calculate GDP NDP PI NI (with numbers , and answer)
Please answer me as fast as possible because tomorrow is my exam
1
Expert's answer
2014-07-03T08:47:51-0400
A) GDP B) NDP C) NI.
All figures are in billions.
1. compensation of employees- $194.2
2. U.S exports of goods and services- $17.8
3. consumption of fixed capital- $11.8
4. governmetn purchases- $59.4
5. taxes on production and imports- $14.4
6. net private domestic investment- $52.1
7. transfer payments- $13.9
8. U.S imports of goods and services- $16.5
9. personal taxes- $40.5
10. net foreign factor income- $2.2
11. personal consumption expenditures- $219.1
12. statistical discrepancy- $0

A)
Expenditure approach:
I = NetInvestment+CapitalConsumption = NetI + CC = 52.1+11.8 = 63.9
GDP = Y = C+I+G+Ex-Im = 219.1+(52.1+11.8) +59.4+17.8-16.5 = 343.7

B)
GDP = NDP + CC
NDP = GDP - CC = 343.7 - 11.8 = 331.9

C)
NI = GNP + SDiscr = GDP+NFI-SDiscr = 331.9+2.2-0 = 334.1
9 is required for disposable income (DI)
7 and 5 are required for personal income (PI)
1 is required to calculate GDP using income method.